OpenStudy (anonymous):

Your piggy bank has a total of 41 coins in it; some are dimes and some are nickels. If you have a total of $3.45, how many nickels do you have?

OpenStudy (sasogeek):

i don't know the value of nickles or dimes in the US currency, sorry

OpenStudy (anonymous):

nickels are .05 dimes are .10

OpenStudy (sasogeek):

let number of nickels = x let number of dimes = y y+x=41 0.05x + 0.1y=3.45 can u solve simultaneous equations?

OpenStudy (anonymous):

15?

OpenStudy (sasogeek):

multiply the second equation by 10, you'll have 0.5x+y=34.5 now subtract your new equation from the first equation x + y = 41 0.5x + y =34.5 ....................... remember that you're subtracting 0.5x +0 = 6.5 this implies that x=6.5/0.5 x=13 if we substitute the value of x into our first equation x+y=41 13+y=41 y=41-13 y=28 so you had 28 dimes and 13 nickels